If you was traveling at 700mph how long would it take to travel 10 miles
Answer by lenny460(1073)

If you was traveling at 700mph how long would it take to travel 10 miles ?
700 / 60 = 11.66 miles per minute
700 / 3600 = 0.194 mile per second
Therefore:
0.194 * 53 = 10.282 miles

Answer:
It will take 53 seconds to travel 10.282 miles










Lennox Obuong
Algebra Tutor
Nairobi, Kenya
Email: obuong3@aol.com